A friend has a 78% average before the final exam for a course. That score includes everything but the final, which counts for 30% of the course grade.
What is the best course grade your friend can earn
What is the minimum score would your friend would need on the final to earn a 75% for the course?

Answer by reviewermath(1029)

A friend has a 78% average before the final exam for a course. That score includes everything but the final, which counts for 30% of the course grade.
What is the best course grade your friend can earn?
best grade = (78%)(0.7) + (100%)(0.3) = 84.6%
What is the minimum score would your friend would need on the final to earn a 75% for the course?
Let x = minimum score
78(0.7) + 0.3x = 75
0.3x = 75 - 78(0.7)
0.3x = 20.4
x = 20.4/0.3
x = 68
Therefore, minimum score is 68%